OpenStudy (kaswanson):

Solve the quadratic equation: -1/2(m-2)(m+1)+5=0

OpenStudy (3mar):

Are you familiar with the quadratic equation solution?

OpenStudy (kaswanson):

I'm familiar with the quadratic equation, but I wasn't sure if there was another way to solve it, or how I would even get it into the form to use the quadratic equation.

OpenStudy (mathmale):

Your "-1/2(m-2)(m+1)+5=0" would be clearer if you were to enclose that fraction, -1/2, inside parentheses: (-1/2)(m-2)(m+1)+5=0. First, multiply both sides of this equation by -2 to clear it of fractions. Second, multiply out the two polynomials on the left side. Third, combine all like terms (e. g., -2+5=3).

OpenStudy (sheriph05):

$-0.5(m-2)(m+1)+5 = 0\\ -0.5(m^2+m-2m-2)+5=0\\ -0.5m^2+0.5m+1+5=0\\ -0.5m^2+0.5m + 6 = 0$\\ complete by using Formula method